OpenStudy (anonymous):

f(x)=-2x^2+2x+3 Find the vertex

OpenStudy (amistre64):

x = -2/-4 = 1/2

OpenStudy (amistre64):

y = -2(1/4) +4 if i see it right

OpenStudy (amistre64):

3.5 ... 7/2.

OpenStudy (amistre64):

v = (1/2 , 7/2) then